# Program evaluation with remotely sensed outcomes

### v2

@@ Abstract
Economists often estimate treatment effects in experiments using remotely sensed variables (RSVs), e.g. satellite images or mobile phone activity, in place of directly measured economic outcomes. A common practice is to use an observational sample to train a predictor of the economic outcome from the RSV, and then to use its predictions as the outcomes in the experiment. We show that this method is biased whenever the RSV is post-outcome, i.e. if variation in the economic outcome causes variation in the RSV. In program evaluation, changes in poverty or environmental quality cause changes in satellite images, but not vice versa. As our main result, we nonparametrically identify the treatment effect by formalizing the intuition that underlies common practice: the conditional distribution of the RSV given the outcome and treatment is stable across the samples. Based on our identifying formula, we find that the efficient representation of RSVs for causal inference requires three predictions rather than one. Valid inference does not require any rate conditions on RSV predictions, justifying the use of complex deep learning algorithms with unknown statistical properties. We re-analyze the effect of an anti-poverty program in India using satellite images.
Keywords: Causal inference, data fusion, treatment-experiments, satellite images
@@ 1 Introduction
While traditional program evaluations rely on surveys to measure impact, important economic outcomes such as living standards and environmental quality may be costly or infeasible to collect. As a consequence, researchers increasingly estimate treatment effects on economic outcomes using remotely sensed variables (RSVs). Examples include night lights as a measure of local economic activity (Chen and Nordhaus,, 2011; Henderson et al.,, 2012; Asher et al.,, 2021); roofing material as a measure of housing quality (Marx et al.,, 2019; Michaels et al.,, 2021; Huang et al.,, 2021); and satellite images as a measure of air pollution (Currie et al.,, 2023), deforestation (Jayachandran et al.,, 2017; Assuncao et al.,, 2023), fires (Jack et al.,, 2025; Balboni et al.,, 2024), flooding (Chen et al.,, 2017; Patel,, 2024), and local poverty (Jean et al.,, 2016; Aiken et al.,, 2022). [^1] Figure 8 illustrates the rapid rise of papers published in economics journals and general interest science journals using RSVs. Of those published in economics journals, we find that 90% use high-dimensional satellite images as RSVs, and 40% use the RSVs as the main outcome in their empirical analysis. See also Burke et al., (2021) and Jack and Walker, (2023) for recent reviews on the empirical uses of RSVs. Our research question is how researchers should rigorously estimate treatment effects from remotely sensed outcomes.
@@
A recurring empirical practice appears in about 50% of the papers in general interest economics journals from 2015-2024 that use remotely sensed outcomes. [^2] We survey AEA journals, Econometrica, Journal of Political Economy, Quarterly Journal of Economics, and Review of Economic Studies. The other 50% use a similar logic, but without an explicit formula for data combination. The common practice is to predict the economic outcome from the RSV, and then to use the predicted outcome in lieu of a true outcome measurement in an experimental sample. Researchers often form such predictions using an auxiliary, observational sample collected in some other context, which contains the RSV and linked outcome measurements. The predictor is typically complex, e.g. a deep learning algorithm with unknown statistical properties.
@@
As motivation, we prove that this intuitive method can lead to positive or negative bias in the estimated treatment effect when the RSV is a post-outcome variable. Using the predicted outcome in lieu of the true outcome implicitly uses the RSV as a surrogate between the treatment and the outcome. However, in many empirical applications using RSVs, the opposite is more plausible: the treatment affects the outcome, and both may affect the RSV. The bias is fundamentally due to this reversal; it is present even without machine learning.
@@
As an example, consider a binary outcome indicating whether a plot of land has been burned (Balboni et al.,, 2024; Jack et al.,, 2025), and an RSV summarizing the color saturation in a satellite image. Fires cause changes in satellite images, but not vice versa; the color saturation of satellite images is post-outcome. A common practice is to predict fires in the experimental sample, using a machine learning algorithm trained on labeled satellite images from an observational sample, and then to compute the difference of predicted outcomes between treated and control units in the experimental sample. Because the RSV is post-outcome, this method’s estimand combines two quantities: the desired effect of the treatment on the outcome, and the correlation between the RSV and the outcome. In the extreme case where the RSV fails to predict the outcome—when an ideal method should report infinite standard errors—this popular method will instead report a precise estimate at zero, regardless of the true treatment effect.
@@
Our main identifying assumptions are jointly testable and lend themselves to simple diagnostics, allowing researchers to assess their plausibility in applications. Of independent interest, we propose a diagnostic to evaluate whether an RSV is relevant enough for an economic outcome to use it for program evaluation.
Our secondary contribution is to characterize the optimal representation of the RSV for inference on the treatment effect. Given our main result, the techniques are standard and hence straightforward to implement; our contribution is to point out the connection between modern remote sensing and classical conditional moments (Chamberlain,, 1987; Newey,, 1993), and to interpret its consequences. We find that three predictions are necessary for semiparametrically efficient downstream causal inference based on RSVs: predictions of the outcome, the treatment, and the sample indicator given the RSV. By contrast, common practice only predicts the outcome given the RSV.
Because modern remote sensing typically involves unstructured data and complex machine learning, we derive valid $n^{-1/2}$ inference without rate conditions, and without complexity restrictions, on RSV-based predictions. Valid inference only requires that (i) a learned RSV representation has some limit; (ii) the limit predicts the outcome of interest, and for this we provide a diagnostic. More precise predictions improve efficiency of program evaluation.
Finally, we conduct a semi-synthetic exercise, calibrated to an existing randomized control trial in India, which we merge with existing satellite images. Following Muralidharan et al., (2016, 2023), we study the effect of Smartcards, a biometrically authenticated payments infrastructure, on village level poverty measures. We use the geographic coordinates of each village to extract nighttime luminosity and high-dimensional embeddings of satellite images. Compared to an unbiased benchmark, estimated from true outcomes in the experiment, our method recovers similar treatment effects with similar precision, despite using remotely sensed variables to compensate for missing experimental outcomes. [^3] Experimental outcomes may be missing for a random half of villages, or for all treated villages. By contrast, common practice can have positive or negative bias for the treatment effect. Conservative cost calculations suggest that using remotely sensed outcomes, instead of directly surveyed outcomes, can recover the treatment effect of interest, while saving about \$$\$3$ million dollars in survey costs.
@@ 1.1 Related work
Compared to various auxiliary variable models in causal inference, we study a different auxiliary variable. Whereas a surrogate is a mediator between the treatment and outcome (Athey et al.,, 2024; Kallus and Mao,, 2024), the RSV is a post-outcome variable in our framework. We show that misusing a post-outcome RSV as a surrogate leads to arbitrary biases for treatment effects. The negative control literature extends the surrogacy framework to deal with unobserved confounding (Ghassami et al.,, 2022; Imbens et al.,, 2024), yet such extensions face the same limitation. See Remark 3.1 for details.
@@
Compared to a vast literature on data combination e.g. Cross and Manski, (2002); Ridder and Moffitt, (2007); Bareinboim and Pearl, (2016); D’Haultfœuille et al., (2025) and nonclassical measurement error e.g. Chen et al., (2011); Schennach, (2020), we place what appears to be a different key assumption. Several influential works handle measurement error in moment condition models by using auxiliary data and assuming that the conditional distribution of the variable of interest, given the imperfect measurement, is stable across samples (Chen et al.,, 2005, 2008; Graham et al.,, 2016), akin to the surrogacy framework. Our key identifying assumption is the opposite: the conditional distribution of the imperfect measurement, given the variable of interest, is stable across samples. This leads to a novel identifying formula.
@@
The main difference between our RSV framework and the prediction powered inference (PPI) framework (Angelopoulos et al.,, 2023; Lu et al.,, 2025; Kluger et al.,, 2025) is also along these lines: the PPI framework uses machine learning predictions as surrogates (Ji et al.,, 2025). Another difference concerns data availability. In our terminology, the PPI approach would require the researcher to observe the treatment, outcome, and RSV for a random subsample of experimental units. The data requirements in other works are similar to those in the PPI literature (Fong and Tyler,, 2021; Allon et al.,, 2023; Gordon et al.,, 2023; Egami et al.,, 2024). By contrast, we allow the researcher to observe no outcomes for any experimental units.
@@
Our results do not require a correctly specified generative model of how treatments and outcomes affect RSVs, which may be prone to mis-specification when the RSV is a satellite image. Several previous works propose methods based on generative modeling (possibly in combination with PPI) that do require correct specification (Gentzkow et al.,, 2019; Alix-Garcia and Millimet,, 2023; Proctor et al.,, 2023; Battaglia et al.,, 2024). Similarly, methods for causal inference on outcomes that are latent concepts require a generative model (Egami et al.,, 2022; Knox et al.,, 2022; Stoetzer et al.,, 2024).
Section 2 formalizes our main assumption: stability of the RSV. Section 3 proves our main result: nonparametric identification. Section 4 characterizes the optimal representation, and provides inference without rate or complexity restrictions. Section 5 shows that our method outperforms common practice in evaluating an anti-poverty program from satellite images.
@@ 2.1 Goal: Identification using remotely sensed outcomes
Within the experimental sample $(S=e)$, we observe pre-treatment covariates $X\in\mathcal{X}$ and a binary treatment $D\in\{0,1\}$. However, the outcome $Y\in\mathcal{Y}$ is missing. [^4] For ease of exposition, we focus on the case where the outcome $Y$ is completely missing in the experimental sample. Remark 3.3 gives the extension where $Y$ is only partially missing in the experimental sample. In its place, we have access to a remotely sensed outcome variable $R\in\mathcal{R}$. We typically think of $R$ as high dimensional (e.g., unstructured data such as satellite images), but it can be low dimensional (e.g., the output of some pre-trained machine learning algorithm). The researcher would like to use the remotely sensed variable (RSV) as an imperfect measurement of the outcome in the experimental sample, without placing parametric assumptions on their relationship.
The causal parameter of interest is the effect of the treatment $D$ on the outcome $Y$ in the experimental sample. Though the outcome $Y$ is unobserved in the experimental sample, we may still define its potential outcomes $Y(d)$ and define our object of interest.
Definition 2.1 (Causal parameter).
@@
Without further assumptions, point identification for this causal parameter is impossible (Horowitz and Manski,, 1995). Even if $R$ can predict $Y$ with great accuracy, if the prediction is at all imperfect, then an assumption is necessary. We therefore place additional structure on the problem, inspired by recent empirical work in environmental and development economics.
A popular practice in environmental and development economics is to use an auxiliary data set of outcomes and RSVs, e.g. of labeled satellite images. We refer to the auxiliary dataset as the observational sample $(S=o)$. For these units, we observe baseline covariates $X$, the outcome $Y$, and the remotely sensed variable $R$. We may or may not observe the treatment $D$. If we do, we denote it by $D\in\{0,1\}$ and refer to this scenario as having “complete” cases. If we do not, or if treatment is deterministic in the observational study, we set $D=0$ for all units in the observational study and refer to this latter scenario as having “incomplete” cases. [^5] Incomplete cases in the observational sample may refer to three scenarios. First, the treatment status may be missing. Second, the treatment status may be present, and all observational units are untreated, hence $D=0$. Third, the treatment status may be present, and all observational units are treated. Redefining treatment gives $\tilde{D}:=1-D=0$. As we discuss below, incomplete cases will require stronger assumptions.
Table 1 summarizes the setting. Each unit is characterized by the random vector $\left(S,X,D,Y(0),Y(1),R\right)$, which we assume to be independent and identically distributed. [^6] Independence is not used to derive our main identification argument, and our framework directly extends to weakly dependent data, as long as a central limit theorem applies. For units in the experimental sample ($S=e$), we observe $(X,D,R)$; for units in the observational sample ($S=o$), we observe $(X,D,Y,R)$ in complete cases or $(X,Y,R)$ in incomplete cases.
@@
Notes: Here, $\checkmark$ denotes the variable is observed. When treatment is missing or deterministic in the observational sample, we encode it as $D=0$ in the observational sample.
Example 2.1 (Environmental impacts).
Consider a randomized experiment that offers cash payments to households in order to incentivize environmental conservation (i.e., “payments for ecosystem services” or PES). Access to PES contracts is often randomized at the village level. We would like to measure whether access to PES contracts $D$ reduces harmful environmental behaviors $Y$, such as deforestation (Jayachandran et al.,, 2017) or crop burning (Jack et al.,, 2025). In a separate observational sample, we link satellite images $R$ to direct measurements $Y$ of deforestation or crop burning e.g. Hansen et al., (2013); Walker et al., (2022). While it is expensive to hire surveyors to record measurements of tree cover or crop management practices in rural areas, it is cheap to collect satellite images. We investigate how to combine these data sources and thereby identify the effect of the PES contracts in the experimental sample. $\blacktriangle$
@@
Figure 1: We illustrate the two samples that we will use to evaluate an anti-poverty program in Andhra Pradesh, India (Muralidharan et al.,, 2023). With experimental units alone and completely missing outcomes, point identification is impossible. Therefore we introduce an auxiliary sample of observational units. See Section 5 for further details.
Example 2.2 (Household poverty).
Consider a randomized experiment evaluating an anti-poverty program, such as an unconditional cash transfer (Egger et al.,, 2022) or biometrically authenticated payment (Muralidharan et al.,, 2023). Treatment is often randomized at the village level. We would like to study the effect of the anti-poverty program $D$ on village-level poverty $Y$. In a separate observational sample, we link satellite images $R$ to census statistics on village-level poverty $Y$. It is well documented that poverty can be predicted from satellite images, with some error e.g. Jean et al., (2016); Rolf et al., (2021). For example, Huang et al., (2021) use deep learning methods to predict household wealth in Kenya from roof quality. While it is expensive to collect poverty measures through in-person surveys in the experimental sample, it is cheap to collect satellite images. We identify the effect of the anti-poverty program in the experimental sample.

### v3

@@ Abstract
Economists often estimate treatment effects in experiments using remotely sensed variables (RSVs), e.g., satellite images or mobile phone activity, in place of directly measured economic outcomes. A common practice is to use an observational sample to train a predictor of the economic outcome from the RSV, and then use these predictions as the outcomes in the experiment. We show that this method is biased whenever the RSV is a post-outcome variable, meaning that variation in the economic outcome causes variation in the RSV. For example, changes in poverty or environmental quality cause changes in satellite images, but not vice versa. As our main result, we nonparametrically identify the treatment effect by formalizing the intuition underlying common practice: the conditional distribution of the RSV given the outcome and treatment is stable across samples. Our identifying formula reveals that efficient inference requires predictions of three quantities from the RSV—the outcome, treatment, and sample indicator—whereas common practice only predicts the outcome. Valid inference does not require any rate conditions on RSV predictions, justifying the use of complex deep learning algorithms with unknown statistical properties. We reanalyze the effect of an anti-poverty program in India using satellite images.
Keywords: Causal inference, data fusion, experiments, satellite images, machine learning.
@@ 1 Introduction
While traditional program evaluations rely on surveys to measure impact, important economic outcomes such as living standards and environmental quality may be costly or infeasible to collect at scale. As a consequence, researchers increasingly estimate treatment effects on economic outcomes using remotely sensed variables (RSVs). Examples include night lights as a measure of local economic activity (Chen and Nordhaus, 2011; Henderson et al., 2012; Asher et al., 2021), roofing material as a measure of housing quality (Marx et al., 2019; Michaels et al., 2021; Huang et al., 2021), mobile phone transactions as a measure of local wealth or consumption (Blumenstock et al., 2015; Aiken et al., 2025), and satellite images as a measure of pollution (Currie et al., 2023), deforestation (Jayachandran et al., 2017; Assuncao et al., 2023), fires (Jack et al., 2025; Balboni et al., 2024), flooding (Chen et al., 2017; Patel, 2024), and local poverty (Jean et al., 2016). [^1] Figure A.1 illustrates the rapid rise of papers published in economics journals and general interest science journals using RSVs. Of those published in economics journals, we find that 90% use high-dimensional satellite images as RSVs, and 40% use the RSVs as the main outcome in their empirical analyses. See also Burke et al. (2021) and Jack and Walker (2023) for recent reviews on the empirical uses of RSVs. Our research question is how researchers should rigorously estimate treatment effects from remotely sensed outcomes.
@@
A recurring empirical practice appears in about 50% of papers in general interest economics journals from 2015-2024 that use remotely sensed outcomes. [^2] We survey AEA journals, Econometrica, Journal of Political Economy, Quarterly Journal of Economics, and Review of Economic Studies. The other 50% of papers use a similar logic, but without an explicit formula for data combination. This common practice predicts the economic outcome from the RSV and then uses the predicted outcome in lieu of a true outcome measurement in an experimental sample. Researchers often form such predictions using an auxiliary, observational sample collected in some other context, which contains the RSV and linked outcome measurements. The predictor is typically complex, e.g., a deep learning algorithm with unknown statistical properties.
@@
We show that this intuitive method can produce arbitrarily biased treatment effect estimates—including flipped signs—when the RSV is a post-outcome variable. Using the predicted outcome in lieu of the true outcome implicitly uses the RSV as a surrogate that mediates between the treatment and the outcome (Prentice, 1989; Athey et al., 2025). However, in many empirical applications using RSVs, the opposite is more plausible: the treatment affects the outcome, and both may affect the RSV. The bias is fundamentally due to this reversal; it is present even without machine learning.
@@
As an example, consider a binary outcome indicating whether a plot of land has been burned (Balboni et al., 2024; Jack et al., 2025) and an RSV summarizing the color saturation in a satellite image. Fires cause changes in satellite images, but not vice versa; the color saturation of satellite images is a post-outcome variable. Common practice predicts fires in the experimental sample using a machine learning algorithm trained on labeled satellite images from an observational sample, and then computes the difference in predicted outcomes between treated and control units in the experimental sample. Because the RSV is post-outcome, this method’s estimand combines two quantities: the desired effect of the treatment on the outcome, and the correlation between the RSV and the outcome. In the extreme case where the RSV fails to predict the outcome—when an ideal method should report infinite standard errors—this popular method will instead report a precise estimate of zero, regardless of the true treatment effect. In general, this method may flip the sign of the true treatment effect.
@@
Our main identifying assumptions are jointly testable and lend themselves to simple diagnostics, allowing researchers to assess their plausibility in applications. We propose a diagnostic to evaluate whether an RSV is sufficiently relevant for an economic outcome to justify its use in program evaluation. Our framework also naturally extends to quasi-experiments, including instrumental variables and difference-in-differences designs, which we present in Appendix J.
Our secondary contribution is to characterize the representation of the RSV that achieves valid, precise, and robust $n^{-1/2}$ inference on the treatment effect. Because modern remote sensing typically involves unstructured data and complex machine learning, we derive valid inference without rate conditions and without complexity restrictions on RSV-based predictions. Valid inference only requires that (i) a learned RSV representation has some limit, and (ii) the limit predicts the outcome of interest, for which we provide a diagnostic. This enables researchers to use complex deep learning algorithms and still conduct valid inference on treatment effects.
Specifically, we build on our main identification result to establish the connection between modern remote sensing and classical conditional moments (Chamberlain, 1987; Newey, 1993). This allows us to derive an expression for a simple representation of the RSV that maximizes precision from the conditional moments we derived. We find that three predictions are necessary for efficient downstream causal inference based on RSVs: predictions of the outcome, the treatment, and the sample indicator given the RSV. By contrast, common practice only predicts the outcome given the RSV. Moreover, $n^{-1/2}$ inference remains valid even if all three predictions are misspecified, which is a strong form of robustness. We provide the remoteoutcome R package to implement our method.
Finally, we conduct a semi-synthetic exercise, calibrated to an existing field experiment in India, which we merge with existing satellite images. Following Muralidharan et al. (2016, 2023), we study the effect of Smartcards, a biometrically authenticated payments infrastructure, on village-level poverty measures. We use the geographic coordinates of each village to extract nighttime luminosity and high-dimensional, pre-trained embeddings of satellite images. Despite using outcomes for only half of the units, we recover the treatment effect of interest with the same precision as an unbiased regression method that has access to outcomes for all units. By contrast, common practice can have positive or negative bias for the treatment effect. Our method enables large savings in survey costs and therefore opens up possibilities for program evaluation in cost-constrained environments.
@@ 1.1 Related Work
Our framework differs from existing approaches to auxiliary variables and data combination along two key dimensions: (i) the causal direction between the auxiliary variable and the outcome, and (ii) the data requirements.
@@
In the surrogacy framework (Prentice, 1989; Athey et al., 2025; Kallus and Mao, 2024), the auxiliary variable (surrogate) is a mediator between the treatment and the outcome, whereas the RSV is a post-outcome variable in our framework. We show that misusing a post-outcome RSV as a surrogate leads to arbitrary bias in treatment effects. The negative control literature extends the surrogacy framework to address unobserved confounding (Ghassami et al., 2022; Imbens et al., 2024), yet such extensions face the same limitation. See Remark 2 for details.
@@
Compared to the vast literature on data combination (e.g., Cross and Manski, 2002; Ridder and Moffitt, 2007; Bareinboim and Pearl, 2016; D’Haultfœuille et al., 2025) and nonclassical measurement error (e.g., Chen et al., 2011; Schennach, 2020), we place what appears to be a different key assumption. Several influential works handle measurement error in moment condition models by using auxiliary data and assuming that the conditional distribution of the variable of interest, given the imperfect measurement, is stable across samples (Chen et al., 2005, 2008; Graham et al., 2016), akin to the surrogacy framework. Our key identifying assumption is the opposite: the conditional distribution of the imperfect measurement, given the variable of interest, is stable across samples. This is natural when the imperfect measurement is caused by the outcome (as with satellite images of poverty or environmental outcomes) rather than merely correlated with it. Other works focus on mismeasured or missing covariates in a sample where the outcome is observed (e.g., Fan et al., 2014; Battaglia et al., 2024; D’Haultfœuille et al., 2025, 2024). By contrast, in our setting, the outcome is missing yet a post-outcome variable is present. Both of these distinctions lead to a novel identifying formula.
@@
The main difference between our RSV framework and the prediction-powered inference (PPI) framework (Angelopoulos et al., 2023; Lu et al., 2025; Kluger et al., 2025) is also along these lines: the PPI framework uses machine learning predictions as surrogates (Ji et al., 2025). Another difference concerns data availability. In our terminology, the PPI approach would require the researcher to observe the treatment, outcome, and RSV for a random subsample of experimental units. The data requirements in other works are similar to those in the PPI literature (Fong and Tyler, 2021; Allon et al., 2023; Gordon et al., 2023; Egami et al., 2024; Carlson and Dell, 2025). By contrast, we allow the researcher to observe no outcomes for any experimental units, enabling inference in settings where PPI-style techniques cannot be applied.
@@
Our results do not require a correctly specified generative model of how treatments and outcomes affect RSVs, which may be prone to misspecification when the RSV is a satellite image or some other type of unstructured data. Several previous works propose methods based on generative modeling that do require correct specification (Gentzkow et al., 2019; Alix-Garcia and Millimet, 2023; Proctor et al., 2023). Similarly, methods for causal inference on outcomes that are latent concepts require a correctly specified generative model (Egami et al., 2022; Knox et al., 2022; Stoetzer et al., 2025).
@@ 2.1 Goal: Identification using Remotely Sensed Outcomes
Within the experimental sample $(S=e)$, we observe pre-treatment covariates $X\in\mathcal{X}$ and a binary treatment $D\in\{0,1\}$. However, the outcome $Y\in\mathcal{Y}$ is missing. [^3] For ease of exposition, we focus on the case where the outcome $Y$ is completely missing in the experimental sample. Remark 4 gives the extension where $Y$ is only partially missing in the experimental sample. In its place, we have access to a remotely sensed outcome variable $R\in\mathcal{R}$. We typically think of $R$ as high-dimensional (e.g., unstructured data such as satellite images), but it could be low-dimensional (e.g., the output of some pre-trained machine learning algorithm). The researcher would like to use the remotely sensed variable (RSV) as an imperfect measurement of the outcome in the experimental sample, without placing parametric assumptions on their relationship.
The causal parameter of interest is the effect of the treatment $D$ on the outcome $Y$ in the experimental sample. Though the outcome $Y$ is unobserved in the experimental sample, we may still define its potential outcomes $Y(d)$ and our parameter of interest. [^4] This definition of potential outcomes has no spillovers. We discuss spillovers in Appendix I.1.
Definition 1 (Causal parameter).
@@
Without further assumptions, point identification for this causal parameter is impossible (Horowitz and Manski, 1995). Even if $R$ can predict $Y$ with great accuracy, if the prediction is at all imperfect, then an assumption is necessary. Therefore, we place additional structure on the problem, inspired by recent empirical work in environmental and development economics.
A popular practice in environmental and development economics is to use an auxiliary dataset of outcomes and RSVs, e.g., of labeled satellite images. We refer to the auxiliary dataset as the observational sample $(S=o)$. For these units, we observe baseline covariates $X$, the outcome $Y$, and the remotely sensed variable $R$. We may or may not observe the treatment $D$. If we do, we denote it by $D\in\{0,1\}$ and refer to this scenario as having “complete” cases. If we do not, or if treatment is deterministic in the observational study, we set $D=0$ for all units in the observational study and refer to this latter scenario as having “incomplete” cases. [^5] Incomplete cases in the observational sample may refer to three scenarios. First, the treatment status may be missing. Second, the treatment status may be present, and all observational units are untreated, hence $D=0$. Third, the treatment status may be present, and all observational units are treated. Relabeling treatment values gives 1D=01-D=0. As we discuss below, incomplete cases will require stronger assumptions.
Table 1 summarizes the setting. Each unit is characterized by the random vector $\left(S,X,D,Y(0),Y(1),R\right)$, which we assume to be independent and identically distributed. [^6] Independence is not used to derive our main identification argument. For units in the experimental sample ($S=e$), we observe $(X,D,R)$; for units in the observational sample ($S=o$), we observe $(X,D,Y,R)$ in complete cases or $(X,Y,R)$ in incomplete cases.
@@
Notes: Here, $\checkmark$ denotes that the variable is observed. When treatment is missing or deterministic in the observational sample, we encode it as $D=0$ in the observational sample.
Example 1 (Environmental impacts).
Consider a randomized experiment that offers cash payments to households in order to incentivize environmental conservation (i.e., “payments for ecosystem services” or PES). Access to PES contracts is often randomized at the village level. We would like to measure whether access to PES contracts $D$ reduces harmful environmental behaviors $Y$, such as deforestation (Jayachandran et al., 2017) or crop burning (Jack et al., 2025). In a separate observational sample, we link satellite images $R$ to direct measurements $Y$ of deforestation or crop burning (e.g., Hansen et al., 2013; Walker et al., 2022). While it is expensive to hire surveyors to record measurements of tree cover or crop management practices in rural areas, it is inexpensive to collect satellite images. We investigate how to combine these data sources and thereby identify the effect of the PES contracts in the experimental sample. $\blacktriangle$
@@
Figure 1: We illustrate the two samples that we will use to evaluate an anti-poverty program in Andhra Pradesh, India (Muralidharan et al., 2023). With experimental units alone and completely missing outcomes, point identification is impossible. Therefore we introduce an auxiliary sample of observational units. See Section 5 for further details.
Example 2 (Household poverty).
Consider a randomized experiment evaluating an anti-poverty program, such as an unconditional cash transfer (Egger et al., 2022) or biometrically authenticated payment (Muralidharan et al., 2023). Treatment is often randomized at the village level. We would like to study the effect of the anti-poverty program $D$ on village-level poverty $Y$. In a separate observational sample, we link satellite images $R$ to census statistics on village-level poverty $Y$. It is well documented that poverty can be predicted from satellite images, with some error (e.g., Jean et al., 2016; Rolf et al., 2021). For example, Huang et al. (2021) use deep learning methods to predict household wealth in Kenya from roof quality. While it is expensive to collect poverty measures through in-person surveys in the experimental sample, it is inexpensive to collect satellite images. We identify the effect of the anti-poverty program in the experimental sample.